A submarine is stranded on the bottom of the ocean with its hatch 23 m below the surface. Calculate the force needed to open the hatch from the inside, given it is circular and 0.42 m in diameter. Air pressure inside the submarine is 1.00 atm


1
Expert's answer
2021-04-19T07:54:33-0400

The pressure of water:


"p=\rho gh."

Force required to open the hatch:


"F=pA=\rho ghA=\rho gh\frac{\pi d^2}{4}.\\\space\\
F=31230\text{ N}."